OpenStudy (anonymous):

solve each system by substitution. 4x-2y=-14 -12x+6y=42

OpenStudy (shane_b):

Can we at least divide everything by 2 first to make it simpler? :)

OpenStudy (anonymous):

?

OpenStudy (shane_b):

Well, all of the terms are divisible by 2 so it's easier to simplify the equations a bit before doing the substitution method. The equations then become: 2x-y=-7 -6x+3y=21

OpenStudy (shane_b):

I would do that first....but you can do it however you like. Substitution method goes like this: 1) Take one of the equations and put x or y in terms of the other variable. Taking the first equation I posted for example: 2x-y=7 so y=2x-7 2) Take this new equation and plug it into one of the equations and solve for x. 3) Once you know what x is, plug it into one of the equations and solve for y.

OpenStudy (shane_b):

4) Check to ensure the solution is valid by plugging x and y into both equations and make sure the equations are still true.